OpenStudy (anonymous):

A cooling tower for a nuclear reactor is to be constructed in the shape of a hyperboloid of one sheet. The diameter at the base is 300 m and the minimum diameter, 500 m above the base, is 220 m. Find an equation for the tower. (Assume the center is at the origin with axis the z-axis and the minimum diameter is at the center.)

OpenStudy (kyosuke):

primero, se sabe que la ecuación de una hiperboloide de una hoja es x^2/a^2+ y^2/b^2+ z^2/c^2=1 ; según el enunciado cuando z=0 ,obtenemos una circunferencia de 110 de radio, por tanto a=110; b=110.luego cuando z=-500 x^2+y^2=150^2 reemplazas en la ecuación y obtienes c=539.31 aproximadamente. con lo hallado armas tu ecuación. suerte.